x + y =5
2x - y =4
-2x +-2y = -10 2x-y =4
2x - y = 4 2x -2 =4

3y -6
__ ___
3 = 3

y = -2 __

then lost - is this right so far and can you show me how to do problem?

Answer by edjones(8007)

x + y =5
2x - y =4
.
x=5-y
2(5-y)-y=4 Replace x with 5-y.
10-2y-y=4
-3y=-6
y=2
x=3
.
Ed